The value of tan 4\(^{\circ}\).tan 43\(^{\circ}\).tan 47 \(^{\circ}\).tan 86 \(^{\circ}\) is
2
3
1
4

Grade:12th pass

1 Answers

Monoxdifly
4 years ago
tan4tan43tan47tan86 = tan4tan43tan(90-43)tan(90-4) = tan4tan43cot43cot4 = tan4cot4tan43cot43 = 1× 1 = 1